# Modelling Lateral Spread in Wire Flat Rolling

### v2

@@ 4.1 Lateral spread
Figure 3: Lateral spread for stainless steel wires with different diameters, $\hat{d}_{0}$, and reduction ratios, $\hat{d}_{0}/(2\hat{h}_{1})$ from different methods; experimental data, empirical equation 2 from kazeminezhad2005width, empirical equation from Kobayashi (3), and the current model. Other parameters used are $(\hat{R},\mu)=(50\,\mathrm{mm},0.25)$.
@@
The results are compared with experimental data, Kobayashi’s empirical equation (3), and empirical equation (2) due to kazeminezhad2005width. The latter equation is particularly relevant because the parameters in (2) were fitted for stainless steel. In the current model, predictions depend on the friction coefficient, which is challenging to determine experimentally. The experimental data were obtained under non-lubricated test conditions, and a friction coefficient of $\mu=0.25$ has been shown to provide the best agreement with experimental data, consistent with the range typically reported for flat rolling of metal wires (lin2008finite; lee2018development). Consequently, this value is used for the results in Figure 3. The current model fails at larger reductions for $\hat{d}_{0}=2.96\ \mathrm{mm}$ and $\hat{d}_{0}=3.96\ \mathrm{mm}$, as indicated by the unfinished lines, due to a singularity in pressure (see equation 24); interestingly, the experiments also encountered difficulties achieving comparable larger reductions during single-pass rolling in these cases, and so experimental data in these cases is also lacking. Where data is available, the model agrees closely with the experimental data for all diameters and reduction ratios. Kazeminezhad Equation (2) depends only on the reduction ratio. While it performs well for smaller reduction ratios, it deviates for larger values, with the deviation appearing to depend on wire diameter (e.g., underestimates for $\hat{d}_{0}=5.96\ \mathrm{mm}$ and overestimates for $\hat{d}_{0}=7.96\ \mathrm{mm}$). Kobayashi’s equation depends on roll radius and wire diameter as well as the reduction ratio and seems to be derived for small wire diameter conditions. It underestimates the lateral spread across all wire diameters, particularly at higher reduction ratios, and is the least reliable for predicting the set of experimental data presented in Figure 3.
@@
Some studies suggest that the friction coefficient has a negligible effect on lateral spread (carlsson1998contact; kazeminezhad2008error), attributing this to the movement of lubricant toward the roll edges, driven by the extremely high contact pressure at the entry point. Similarly, the empirical equations (2) and (3) exhibit no dependence on friction. To examine the impact of friction, the ratio of lateral spread to the initial wire diameter is plotted in Figure 4 for two different friction coefficients. In the absence of experimental data under the lubricated condition, the FE results from vallellano2008analysis are utilised. The data reported by vallellano2008analysis corresponds to a wire with a 5 $\mathrm{mm}$ diameter, rolled using 75 $\mathrm{mm}$ radius rolls, with no applied forward or backward tension. The yield stress, $\hat{Y}$, is assumed to be constant and equal to 385 $\mathrm{MPa}$ the same as the average value in FE simulations. vallellano2008analysis used the Tresca friction law, $\hat{\tau}=m\hat{\kappa}$ with the coefficient of $m=0.25$ to generate their FE results. Generally, the relationship $\mu<m/\sqrt{3}$ is used essentially for a simple elastoplastic model (zhang2016relationship), where the upper limit, $\mu=0.14$, is used to transfer $m$ into the Coulomb friction coefficient, $\mu$. The data represented by the black line with circles corresponds to the experimental measurements for $\hat{d}_{0}=5.96~\mathrm{mm}$ without lubrication ($\mu=0.25$), while the counterpart for $\hat{d}_{0}=5~\mathrm{mm}$ is derived from FE simulations using $m=0.25$ (approximated by $\mu=0.14$). To minimise the influence of different wire diameters in this comparison, wires with relatively similar diameters were selected, and the results were scaled by the initial diameter.
@@
Figure 4: Effect of friction coefficient on lateral spread. Results are plotted from different methods; experimental data, FE simulations (vallellano2008analysis), empirical equation 2 from kazeminezhad2005width, empirical equation from Kobayashi (3), and the current model.
@@ 4.3 Roll pressure
Figure 6: Roll pressure from different methods; plane-strain slab method (erfanian2025), current plane-stress method, and FE simulations (vallellano2008analysis) and kazeminezhad2006pressure slab method. Plotted for different reduction ratios against contact angle $\phi$. Other parameters used are $(\hat{\kappa},\hat{d}_{0},\hat{R},\mu)=(385/\sqrt{3}\,\mathrm{MPa},5\,\mathrm{mm},75\,\mathrm{mm},0.14)$.
Results are compared with FE simulations from vallellano2008analysis, the slab model from kazeminezhad2006pressure and the plane-strain slab model for the rolling of thin sheet in erfanian2025. The same parameters as those used in vallellano2008analysis and described in the previous section are applied. The results for plane stress and plane strain are calculated from point A in Figure 2.
@@
Figure 7: A diagram of the deforming wire, seen from above (top figure) and from side (bottom figure). For symbols, refer to the main text. Adapted from Figure 3 of carlsson1998contact.
which schematically shows how the wire looks as seen from the side and above while rolling. The wire enters between rolls from the right-hand side with a circular cross-section and exits from the left with an almost rectangle cross-section, therefore, the contact surface forms rather like half of an ellipse. As explained by carlsson1998contact when the material starts to deform, surrounding parts of the wire are still in the elastic range and will therefore resist deformation. The situation may be compared to that of an indentation. As a result, a normal pressure will build up, resulting in a sharp rise in contact pressure soon after the entrance. As deformation continues, larger parts of the wire deform, reducing resistance from the remaining elastic regions, which in turn results in a pressure drop until point A. At this point, the material begins to flow laterally, and a typical friction hill develops, with a pressure peak forming between this point and the roll exit (carlsson1998contact).
From the results, it can be seen that the current model correctly predicts the location of point A by assuming that the area of the cross-section at this point is the same as that of the wire before rolling. Both the current plane-stress model and kazeminezhad2006pressure correctly predict the roll pressure at point A, for larger reductions. From point A to the roll exit, the FE results lie between the plane-strain and plane-stress predictions, reflecting the presence of 3D material flow within the roll gap. Consequently, the results presented by kazeminezhad2006pressure show better agreement with the FE data. Yet, despite the model underestimates the pressure, it consistently provides accurate predictions of lateral spread. This may be attributed to the importance of the stress ratio $s_{zz}/s_{yy}$ — rather than the absolute magnitudes — in governing the lateral deformation of the wire. The model appears to capture this ratio correctly, despite discrepancies in individual stress components. Another notable difference between the FE results and the perfect plastic models in Figure 6 is that the roll pressure starts at zero at the entrance and returns to zero at the exit. This behaviour arises from the inclusion of elasticity in the FE simulations, which is absent in the perfect plastic models.
@@ 6 Conclusion
Nonetheless, having formulated stress and strain equations in all directions lays the foundation for a more general model. 3D numerical studies conducted using Abaqus software by carlsson1998contact and vallellano2008analysis show that the contact pressure distribution is more complex than the traditional pressure hill profile. Relaxing the assumption of a small rate of change in width in the current model would allow the problem to be modelled under a general stress state. Also, solving the higher order terms would enable accounting for the curvature of the edges.
Cold-rolled wire often exhibits a strong texture in the axial direction, while the texture in the cross-section can be quite different (carlsson1998contact). The von Mises function used to describe the yielding condition could be replaced with an anisotropic yield criterion (e.g., Hill’s 1948 yield criterion) with known anisotropy coefficients. In such a case, the solution structure would largely remain unchanged, except for the incorporation of weighted stresses. However, this scenario is of limited scientific significance because it is the rolling process itself which induces anisotropy. To accurately capture the anisotropy effect, it would be necessary for the anisotropy to evolve throughout the process. This would require a model describing how deformation induces anisotropy within the material, thereby adding an additional layer of complexity to the modelling process and is currently work in progress. Additionally, further studies are required to explore potential correlations between the challenges faced in experiments when achieving certain reductions in a single pass and the pressure singularity predicted by the current model.
@@ Appendix A Effect of initial cross-sectional shape
All results here have assumed that when a cylindrical wire is rolled, it first transforms into a square wire with aspect ratio $\hat{b}_{\!\>\!\mathrm{A}}/\hat{h}_{\!\>\!\mathrm{A}}=a=1$, as this preserves the symmetry of the cylindrical wire. In this appendix, we further investigate the effects of this assumption.
Figure 10: Effect of the assumed cross-sectional shape at location A on the predicted lateral spread. Different shapes corresponding to different values of $a=\hat{b}_{A}/\hat{h}_{A}$ are illustrated in the left panel.
Figure 10, computed as described in section 2.4 but for $b_{A}=a$, shows similar results to figure 4 but details the effect of varying $a$ on the results of our model. Figure 10 justifies that the choice of $a=1$ not only preserves symmetry but also gives the most accurate results compared with experimental results. Similar results were obtained for all experimental wire diameters, although Figure 10 plots the experimental results for the $\hat{d}_{0}=7.96\,\mathrm{mm}$ wire as these showed the greatest variation with $a$.

### v3

@@ 4.1 Lateral spread
Figure 3: Lateral spread for stainless steel wires with different diameters, $\hat{d}_{0}$, and reduction ratios, $\hat{d}_{0}/(2\hat{h}_{1})$ from different methods; experimental data, empirical equation 2 from Kazeminezhad and Karimi Taheri (2005), empirical equation from Kobayashi (3), and the current model. Other parameters used are $(\hat{R},\mu)=(50\,\mathrm{mm},0.25)$.
@@
The results are compared with experimental data, Kobayashi’s empirical equation (3), and empirical equation (2) due to Kazeminezhad and Karimi Taheri (2005). The latter equation is particularly relevant because the parameters in (2) were fitted for stainless steel. In the current model, predictions depend on the friction coefficient, which is challenging to determine experimentally. The experimental data were obtained under non-lubricated test conditions, and a friction coefficient of $\mu=0.25$ has been shown to provide the best agreement with experimental data, consistent with the range typically reported for flat rolling of metal wires (Lin et al., 2008; Lee et al., 2018). Consequently, this value is used for the results in Figure 3. The current model fails at larger reductions for $\hat{d}_{0}=2.96\ \mathrm{mm}$ and $\hat{d}_{0}=3.96\ \mathrm{mm}$, as indicated by the unfinished lines, due to a singularity in pressure (see equation 24); interestingly, the experiments also encountered difficulties achieving comparable larger reductions during single-pass rolling in these cases, and so experimental data in these cases is also lacking. Where data is available, the model agrees closely with the experimental data for all diameters and reduction ratios. Kazeminezhad Equation (2) depends only on the reduction ratio. While it performs well for smaller reduction ratios, it deviates for larger values, with the deviation appearing to depend on wire diameter (e.g., underestimates for $\hat{d}_{0}=5.96\ \mathrm{mm}$ and overestimates for $\hat{d}_{0}=7.96\ \mathrm{mm}$). Kobayashi’s equation depends on roll radius and wire diameter as well as the reduction ratio and seems to be derived for small wire diameter conditions. It underestimates the lateral spread across all wire diameters, particularly at higher reduction ratios, and is the least reliable for predicting the set of experimental data presented in Figure 3.
@@
Some studies suggest that the friction coefficient has a negligible effect on lateral spread (Carlsson, 1998; Kazeminezhad et al., 2008), attributing this to the movement of lubricant toward the roll edges, driven by the extremely high contact pressure at the entry point. Similarly, the empirical equations (2) and (3) exhibit no dependence on friction. To examine the impact of friction, the ratio of lateral spread to the initial wire diameter is plotted in Figure 4 for two different friction coefficients. In the absence of experimental data under the lubricated condition, the FE results from Vallellano et al. (2008) are utilised. The data reported by Vallellano et al. (2008) corresponds to a wire with a 5 $\mathrm{mm}$ diameter, rolled using 75 $\mathrm{mm}$ radius rolls, with no applied forward or backward tension. The yield stress, $\hat{Y}$, is assumed to be constant and equal to 385 $\mathrm{MPa}$ the same as the average value in FE simulations. Vallellano et al. (2008) used the Tresca friction law, $\hat{\tau}=m\hat{\kappa}$ with the coefficient of $m=0.25$ to generate their FE results. Generally, the relationship $\mu<m/\sqrt{3}$ is used essentially for a simple elastoplastic model (Zhang and Ou, 2016), where the upper limit, $\mu=0.14$, is used to transfer $m$ into the Coulomb friction coefficient, $\mu$. The data represented by the black line with circles corresponds to the experimental measurements for $\hat{d}_{0}=5.96~\mathrm{mm}$ without lubrication ($\mu=0.25$), while the counterpart for $\hat{d}_{0}=5~\mathrm{mm}$ is derived from FE simulations using $m=0.25$ (approximated by $\mu=0.14$). To minimise the influence of different wire diameters in this comparison, wires with relatively similar diameters were selected, and the results were scaled by the initial diameter.
@@
Figure 4: Effect of friction coefficient on lateral spread. Results are plotted from different methods; experimental data, FE simulations (Vallellano et al., 2008), empirical equation 2 from Kazeminezhad and Karimi Taheri (2005), empirical equation from Kobayashi (3), and the current model.
@@ 4.3 Roll pressure
Figure 6: Roll pressure from different methods; plane-strain slab method (Erfanian et al., 2025), current plane-stress method, and FE simulations (Vallellano et al., 2008) and Kazeminezhad and Karimi Taheri (2006) slab method. Plotted for different reduction ratios against contact angle $\phi$. Other parameters used are $(\hat{\kappa},\hat{d}_{0},\hat{R},\mu)=(385/\sqrt{3}\,\mathrm{MPa},5\,\mathrm{mm},75\,\mathrm{mm},0.14)$.
Results are compared with FE simulations from Vallellano et al. (2008), the slab model from Kazeminezhad and Karimi Taheri (2006) and the plane-strain slab model for the rolling of thin sheet in Erfanian et al. (2025). The same parameters as those used in Vallellano et al. (2008) and described in the previous section are applied. The results for plane stress and plane strain are calculated from point A in Figure 2.
@@
Figure 7: A diagram of the deforming wire, seen from above (top figure) and from side (bottom figure). For symbols, refer to the main text. Adapted from Figure 3 of Carlsson (1998).
which schematically shows how the wire looks as seen from the side and above while rolling. The wire enters between rolls from the right-hand side with a circular cross-section and exits from the left with an almost rectangle cross-section, therefore, the contact surface forms rather like half of an ellipse. As explained by Carlsson (1998) when the material starts to deform, surrounding parts of the wire are still in the elastic range and will therefore resist deformation. The situation may be compared to that of an indentation. As a result, a normal pressure will build up, resulting in a sharp rise in contact pressure soon after the entrance. As deformation continues, larger parts of the wire deform, reducing resistance from the remaining elastic regions, which in turn results in a pressure drop until point A. At this point, the material begins to flow laterally, and a typical friction hill develops, with a pressure peak forming between this point and the roll exit (Carlsson, 1998).
From the results, it can be seen that the current model correctly predicts the location of point A by assuming that the area of the cross-section at this point is the same as that of the wire before rolling. Both the current plane-stress model and Kazeminezhad and Karimi Taheri (2006) correctly predict the roll pressure at point A, for larger reductions. From point A to the roll exit, the FE results lie between the plane-strain and plane-stress predictions, reflecting the presence of 3D material flow within the roll gap. Consequently, the results presented by Kazeminezhad and Karimi Taheri (2006) show better agreement with the FE data. Yet, despite the model underestimates the pressure, it consistently provides accurate predictions of lateral spread. This may be attributed to the importance of the stress ratio $s_{zz}/s_{yy}$ — rather than the absolute magnitudes — in governing the lateral deformation of the wire. The model appears to capture this ratio correctly, despite discrepancies in individual stress components. Another notable difference between the FE results and the perfect plastic models in Figure 6 is that the roll pressure starts at zero at the entrance and returns to zero at the exit. This behaviour arises from the inclusion of elasticity in the FE simulations, which is absent in the perfect plastic models.
@@ 6 Conclusion
Figure 10: Effect of the assumed cross-sectional shape at location A on the predicted lateral spread. Different shapes corresponding to different values of $a=\hat{b}_{\!\>\!\mathrm{A}}/\hat{h}_{\!\>\!\mathrm{A}}$ are illustrated in the left panel.
@@
Nonetheless, having formulated stress and strain equations in all directions lays the foundation for a more general model. 3D numerical studies conducted using Abaqus software by Carlsson (1998) and Vallellano et al. (2008) show that the contact pressure distribution is more complex than the traditional pressure hill profile. Relaxing the assumption of a small rate of change in width in the current model would allow the problem to be modelled under a general stress state. Also, solving the higher order terms would enable accounting for the curvature of the edges.
Cold-rolled wire often exhibits a strong texture in the axial direction, while the texture in the cross-section can be quite different (Carlsson, 1998). The von Mises function used to describe the yielding condition could be replaced with an anisotropic yield criterion (e.g., Hill’s 1948 yield criterion) with known anisotropy coefficients. In such a case, the solution structure would largely remain unchanged, except for the incorporation of weighted stresses. However, this scenario is of limited scientific significance because it is the rolling process itself which induces anisotropy. To accurately capture the anisotropy effect, it would be necessary for the anisotropy to evolve throughout the process. This would require a model describing how deformation induces anisotropy within the material, thereby adding an additional layer of complexity to the modelling process and is currently work in progress. Additionally, further studies are required to explore potential correlations between the challenges faced in experiments when achieving certain reductions in a single pass and the pressure singularity predicted by the current model.
@@ Appendix A Effect of initial cross-sectional shape
All results here have assumed that when a cylindrical wire is rolled, it first transforms into a square wire with aspect ratio $\hat{b}_{\!\>\!\mathrm{A}}/\hat{h}_{\!\>\!\mathrm{A}}=a=1$, as this preserves the symmetry of the cylindrical wire. In this appendix, we further investigate the effects of this assumption. Figure 10, computed as described in section 2.4 but for $b_{\!\>\!\mathrm{A}}=a$, shows similar results to figure 4 but details the effect of varying $a$ on the results of our model. Figure 10 justifies that the choice of $a=1$ not only preserves symmetry but also gives the most accurate results compared with experimental results. Similar results were obtained for all experimental wire diameters, although Figure 10 plots the experimental results for the $\hat{d}_{0}=7.96\,\mathrm{mm}$ wire as these showed the greatest variation with $a$.